OpenStudy (campayne):

what is the augmented matrix y - 3 = 2x 3x = -y

jimthompson5910 (jim_thompson5910):

I'll show you how to do a similar example. Let's say we had this system y - 5 = 9x 4x = -2y ------------------------------------------------------------------------------------------ The goal is to get the two equations into `Ax+By = C` form. To do that for the first equation, we subtract 9x from both sides and add 5 to both sides y - 5 = 9x y - 5-9x = 9x-9x y - 5-9x = 0 y - 5-9x+5 = 0+5 -9x+y = 5 Now onto the second equation 4x = -2y. Let's add 2y to both sides to get 4x = -2y 4x+2y = -2y+2y 4x+2y = 0 ------------------------------------------------------------------------------------------ So we have this new equivalent system -9x+y = 5 4x+2y = 0 which is the same as this system -9x+1y = 5 4x+2y = 0 The system above is in the form Ax+By = C Dx+Ey = F where... A = -9 B = 1 C = 5 D = 4 E = 2 F = 0 Now just plug all this into the augmented matrix below \[\Large \left[\begin{array}{cc|c}A & B & C\\D & E & F\end{array}\right]\] to get \[\Large \left[\begin{array}{cc|c}-9 & 1 & 5\\4 & 2 & 0\end{array}\right]\] and that's all there is to it. Hopefully this example is enough to get you pointed in the right direction to answer your question.